OpenStudy (anonymous):

What is the area under the curves of y=3, 2y=5(x)^1/2 and 2y+x=6

OpenStudy (anonymous):

is there a picture attached?

OpenStudy (anonymous):

Ah well it's just that graphing this is going to be a pain.. So let's see what we can do

OpenStudy (anonymous):

no but I can attach one

OpenStudy (anonymous):

if u can, please do so haha

OpenStudy (anonymous):

k here it is

OpenStudy (anonymous):

omg lol more of these webworks... i'm so sick of them haha

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

nah it's fine.. actually give the wolframalpha link, they also list the intercepts. it's easier.

OpenStudy (anonymous):

all right great thanks. Here we go..

OpenStudy (anonymous):

So here I'm going to split the graph in two at the point where x+2y = 6 and 2y = 5sqrt(x) intersect

OpenStudy (anonymous):

okay we did that

OpenStudy (anonymous):

That's the upper limit of our first integral

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Sorry my browser's having trouble, anyway the upper intersect was 1. Now the lower intersect, which occurs when y =3 and 2y+x=6 intersect.

OpenStudy (anonymous):

6+x=6, x = 0. So the first half of that area is: Integral (upper curve - lower curve, 0,1)

OpenStudy (anonymous):

how did you get one I got 1.36 doesnt it turn into a quadratic

OpenStudy (anonymous):

x+5sqrt(x) = 6? Well I mean take a look, what number plus 5 times it's square root is 6?

OpenStudy (anonymous):

okay i get it i never thought of it like that

OpenStudy (anonymous):

Keep in mind, we're still calculating the first half. So now, the upper curve minus the lower curve (well they're lines, but whatever) is: 3 - (6-x)/2

OpenStudy (anonymous):

So so far, the first half of our area is: Integral (3 - (6-x)/2, 0,1)

OpenStudy (anonymous):

Now let's find the 2nd half: This time the lower limit is 1, because we are starting at one, and the upper limit is the intersection of y=3 and y = 5/2sqrt(x)

OpenStudy (anonymous):

(5/2)*sqrt(x) = 3 sqrt(x) = 6/5 => x = 36/25

OpenStudy (anonymous):

So 2nd integral is again, upper curve minus the lower curve from 1 to 36/25, or Integral (3 - (5/2)*sqrt(x), 1, 36/25)

OpenStudy (anonymous):

Calculate each integral and then add them to get 107/300 (lol I cheated, used wolf to compute)

OpenStudy (anonymous):

so it is .356666667

OpenStudy (anonymous):

Yea that looks right.

OpenStudy (anonymous):

k i will try it and tell you if it works

OpenStudy (anonymous):

it works thanks sooooo much you have no idea

OpenStudy (anonymous):

np.